OpenStudy (anonymous):

Which is the solution of the equation below, rounded to the nearest tenth? 1.6x – 3.38 = 5.81x + 4.67 A. –0.2 B. –2.0 C. –0.3 D. –1.9

OpenStudy (texaschic101):

1.6x - 3.38 = 5.81x + 4.67 (subtract 5.81x from both sides) 1.6x - 5.81x - 3.38 = 4.67 (add 3.38 to both sides) 1.6x - 5.81x = 4.67 + 3.38 (combine like terms) -4.21x = 8.05 (divide both sides by -4.21) x = 8.05/-4.21 x = - 1.91 rounds to - 1.9

OpenStudy (anonymous):

thanks :) @texaschic101

OpenStudy (texaschic101):

no problem :)